Find all real and imaginary solutions
w^2 = -225
Could someone please help me with this, I am drawing a blank!

Answer by stanbon(75887)

Find all real and imaginary solutions
---
Factors of a^2+b^2 are (a+bi) and (a-bi)
-----------------------
Your problem:
w^2 = -225
w^2+225 = 0
---
w^2+15^2 = 0
---
(w+15i)(w-15i) = 0
w = -15i or w = +15i
===================
Cheers,
Stan H.